XxShado:

Find the product of 3 1/5 and 7 1/4. Express your answer in simplest form. Answers: 21 1/20 22 10 1/9 23 1/5

XxShado:

ok please help

mymommy:

hope this helps

crispyrat:

ok so 3 1/5=(3*5+1)/5=16/5 and then 7 1/4 is (7*4+1)/4=29/4 Thus (16/5)*(29/4 )is 494/20 which is 23 1/5